OpenStudy (anonymous):

4. Solve 3x + 7 = -5 for x. (Points : 4) -4 -2 2 4 5. Solve -3x + 2 = 23 for x. (Points : 4) 3 7 -3 -7 6. Solve 2x - 6 = 10 for x. (Points : 4) 4 6 8 10

OpenStudy (ja1):

3x + 7 = -5 - 7 - 7 3x = -12 /3 /3 x = -4

OpenStudy (ja1):

-3x + 2 = 23 - 2 - 2 -3x = 21 /-3 /-3 x = -7

OpenStudy (ja1):

2x - 6 = 10 + 6 + 6 2x = 16 /2 /2 x = 8

OpenStudy (ja1):

I hope your getting this @jayfer

OpenStudy (anonymous):

you're always just working on isolating the x value. So you just want to get the x on one side, and the other values on the other side. Simplify till you get the value. Just follow what @JA1 wrote :)